OpenStudy (anonymous):

what is the answer to 2x3-10x2=0 and w2-3w=0

OpenStudy (anonymous):

2x^3-10x^2=0 2x^2(x-5)=0 2x^2=0 x-5=0 x=0 x=5

OpenStudy (anonymous):

w^2-3w=0 w(w-3)=0 w=0 w-3=0 w=3

OpenStudy (anonymous):

thanks